--------------------------------------------------------------------------------
Investing in stocks.. US small company stocks have returned an average of 14.9% annually for the last 50 years. Use the present value formula from the previous exercise to find the amount invested today in small company stocks that would be worth $1 million in 50 years, assuming that small company stocks continue to return 14.9% annually for the next 50 years.
The previous question (present value) P=A(1+r)^-n the present value P that will amount to A dollars in n years with interest compounded annually at an anual interest rate r, is given by P=A(1+r)^-n
Find the present value that will amount to $50,000 in 20 years at 8% compounded annually.
Thanks for any light you can shed on this. I don't know where to begin and I need to show how I did it. Sheryl
: --------------------------------------------------------------------------------
Investing in stocks.. US small company stocks have returned an average of 14.9% annually for the last 50 years. Use the present value formula from the previous exercise to find the amount invested today in small company stocks that would be worth $1 million in 50 years, assuming that small company stocks continue to return 14.9% annually for the next 50 years.
The previous question (present value) P=A(1+r)^-n the present value P that will amount to A dollars in n years with interest compounded annually at an anual interest rate r, is given by P=A(1+r)^-n
Find the present value that will amount to $50,000 in 20 years at 8% compounded annually.
Thanks for any light you can shed on this. I don't know where to begin and I need to show how I did it. Sheryl

Answer by stanbon(18327)

The previous question (present value) P=A(1+r)^-n the present value P that will amount to A dollars in n years with interest compounded annually at an anual interest rate r, is given by P=A(1+r)^-n
Find the present value that will amount to $50,000 in 20 years at 8% compounded annually.
---------------------
P = 50000(1+0.08)^(-20)
P = 50000(1.08)^(-20)
P = $10727.41
==================
Cheers,
Stan H.